# Improved Patch Denoising Diffusion Probabilistic Models for Magnetic Resonance Fingerprinting

### v1

@@ 1 Introduction
Quantifying the intrinsic nuclear magnetic resonance (NMR) properties of tissues [1] provides a powerful tool for tissue characterization and monitoring of pathological changes. However, conventional quantitative MRI (qMRI) techniques face challenges due to long acquisition times, limiting their clinical utility. This is because qMRI often requires sampling a long sequence (timeseries) of magnetization images to capture tissue properties, making it difficult to acquire the full spatiotemporal volume in a reasonable scan time. Magnetic Resonance Fingerprinting (MRF) [2] and other transient-state imaging techniques [3, 4, 5] were introduced to address this problem. To reduce scan time, MRF uses short-length acquisition sequences to encode multiple tissue properties simultaneously and applies compressed sensing to subsample only a fraction of the spatio-temporal k-space data. However, faster scans lead to challenges in image reconstruction, including aliasing artifacts from k-space undersampling and limited tissue property information due to the truncated acquisition sequences. Effective image reconstruction algorithms are needed to tackle these challenges and improve the accuracy and precision of tissue parameter estimation.
@@
Early MRF imaging algorithms [2, 6] used gridding reconstruction (non-Cartesian Fourier back projection) to convert undersampled k-space data into image timeseries. Tissue properties were then estimated through dictionary matching, where voxel timesignals were compared to a precomputed dictionary of theoretical tissue fingerprints. However, the accuracy of these methods was limited by aliasing artifacts that were not addressed during the image reconstruction step. While later compressed sensing algorithms [7, 8, 9, 10, 11] introduced various sparsity and/or low-rank priors to regularize reconstruction and reduce artifacts, the most effective MRF algorithms to date are data driven, based on deep learning [12, 13, 14, 15, 16, 17, 18]. These methods use some form of convolutional neural networks (CNNs) to directly learn spatio-temporal priors from the data, allowing them to more accurately restore (de-alias) MRF image timeseries or reconstruct quantitative maps.
@@
Recently denoising diffusion probabilistic models (DDPM) [19, 20, 21, 22] have gained significant attention due to their remarkable performance on image generation [23, 24, 25] and image restoration tasks [26, 27, 28, 29, 30]. Built on a hierarchy of denoising autoencoders, DDPMs are trained to iteratively refine image details by denoising and reversing a diffusion process. While their applications in medical imaging are emerging (e.g. in segmentation [31, 32], image-to-image translation [33, 34], and image reconstruction [35, 36, 37]), their potential for MRF imaging is yet unexplored. Compared to CNNs, the current state-of-the-art for MRF reconstruction, DDPMs offer several advantages. DDPMs’ iterative refinement allows for stage-wise, progressive restoration of image details from coarse to fine scales, leading to more precise recovery compared to the single-pass structure of CNNs. They also incorporate advanced mechanisms like self-attention [38] or Transformers [39], which more effectively model long-range pixel dependencies. Additionally, while CNNs treat image restoration as a deterministic mapping from a corrupted to a restored image, DDPMs operate within a probabilistic framework, allowing them to learn the distribution of possible solutions for a given task and generate multiple samples i.e. a useful feature for capturing uncertainties in the reconstructed images.
@@
This paper presents, to the best of our knowledge, the first diffusion-based model for MRF image reconstruction. We train a conditional DDPM model to generate restored MRF image timeseries from initial gridding reconstructions, which suffer from aliasing artifacts due to accelerated MRF acquisitions. Evaluated on in-vivo brain scans data, we show that the proposed approach outperforms established deep learning and compressed sensing MRF algorithms in terms of image reconstruction and quantitative mapping accuracies. Additionally, the probabilistic nature of DDPM allows us to generate multiple sample reconstructions given an input and produce a variance map to visualize reconstruction uncertainties. To address the computational challenges associated with training diffusion models on high-dimensional MRF data, we employed a subspace dimensionality reduction approach from [6] to temporally compress the MRF image timeseries, enabling the diffusion process to operate within a reduced latent space. We further implemented a patch-based processing pipeline to efficiently handle the spatial dimensions of data, enabling training and reconstruction on modest GPUs with limited memory and reasonable runtimes.
@@ 2.1 The MRF Reconstruction Problem
The MRF image reconstruction is an inverse problem:
@@
where, $\textbf{y}\in\mathbb{C}^{c\times m\times l}$ are the undersampled k-space measurements from $m$ spatial frequency locations across $c$ receiver coil channels, and $l$ timeframes. The linear acquisition operator $\underline{A}:\mathbb{C}^{n\times l}\rightarrow\mathbb{C}^{c\times m\times l}$ includes coil sensitivities and the undersampled spatial Fourier operations at each timeframe. Given $\mathbf{y}$, the aim is to reconstruct the timeseries of magnetization images (TSMI) $\mathbf{\underline{x}}\in\mathbb{C}^{n\times l}$ that contains the tissue-encoding signals of length $l$ for $n$ voxels. The TSMI and the desired tissue parameter maps (Q-Maps), here the spatial maps of T1 and T2 relaxations times, are voxel-wise related through the non-linear Bloch response model $\mathcal{B}:\mathbb{R}^{2}\rightarrow\mathbb{C}^{l}$, scaled by proton density $\boldsymbol{\rho}$. The MRF dictionary D is a matrix that contains precomputed Bloch responses for a discretized grid of (T1, T2) values, and it acts as a lookup table linking tissue properties to corresponding magnetization time signals or fingerprints. Once the TSMI $\mathbf{\underline{x}}$ is reconstructed, a voxel-wise dictionary matching process is applied to identify the T1 and T2 maps [2]. In this way, the dictionary matching step effectively inverts the non-linear Bloch response model. McGivney et al. [6] proposed a low-rank SVD decomposition of the MRF dictionary, approximating $\textbf{D}\approx\textbf{\text@underline{D}}\textbf{V}^{H}$, where D is the dimensionally-reduced dictionary in an $s$-dimensional subspace ($s\ll l$) defined by the orthonormal basis $\textbf{V}\in\mathbb{C}^{l\times s}$. They also show that TSMI can be factorized in this basis as $\mathbf{\underline{x}}\approx\mathbf{\underline{x}}\textbf{VV}^{H}=\mathbf{x}\textbf{V}^{H}$, where $\mathbf{x}=\underline{x}\textbf{V}$ represents the dimensionally-reduced TSMI. With this reduction, our reconstruction problem becomes estimating a dimension-reduced TSMI from $\mathbf{y}\approx A\mathbf{x}$, where $A=\underline{A}\circ\textbf{V}$. Unless otherwise specified, this dimension reduction is consistently applied throughout our work.
The MRF acquisition process can be greatly shortened by combining k-space undersampling with a reduction in the number of measured timeframes $l$. However, this approach introduces severe aliasing artifacts and increased noise sensitivity due to the limited tissue property information in the reconstructed TSMI — issues we aim to address in this work.
@@ 2.2 Denoising Diffusion Probabilistic Models for Image Reconstruction
Denoising Diffusion Probabilistic Models (DDPM) [19, 20] describe a Markovian process in which a sample image from the data distribution $q(\mathbf{x}_{0})$ is gradually corrupted by adding Gaussian noise over $t=1,\ldots,T$ diffusion steps, generating latent variables $\mathbf{x}_{t}$ with increasing noise levels $\beta_{t}\in(0,1)$. For a sufficiently large $T$, the variable $\mathbf{x}_{T}$ approximates an isotropic Gaussian distribution, effectively becoming pure noise. The stepwise forward process is defined as:
$q(\mathbf{x}_{t},|\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\textbf{Id}),$ (2)
@@
without having to progress through all previous diffusion steps. The DDPM models train a neural network $\epsilon_{\theta}$ to estimate the noise $\epsilon$ at each diffusion step, with a learning objective to minimize the loss:
@@ 3.1 Preprocessing
In the preprocessing step, we start with k-space measurements $\mathbf{y}_{\text{ref}}$ from extended reference MRF acquisitions such as [3] that can be used to generate low-aliasing quantitative maps via compressed sensing reconstruction methods like LRTV [40]. From these quantitative maps we then produce high-quality (low-aliasing) reference TSMIs $\mathbf{x}_{\text{ref}}$ using the Bloch response model in Eq. (1). Next, we subsample $\mathbf{y}_{\text{ref}}$ by truncating it along the time dimension to obtain k-space data $\mathbf{y}$, representing retrospectively faster acquisitions. Reconstructions of these accelerated acquisitions using a basic gridding algorithm $\mathbf{x}_{\text{grid}}=A^{H}\mathbf{y}$, or a compressed sensing method contain severe aliasing artifacts.
@@ 3.2 Training
The training process is outlined in Algorithm 1. We choose a denoising model $\epsilon_{\theta}$ from [21]. The model is trained on TSMI patches $\{\mathbf{x}_{c}^{(i)}=\text{Crop}(\textbf{P}_{i},\mathbf{x}_{c})$, $\mathbf{x}_{0}^{(i)}=\text{Crop}(\textbf{P}_{i},\mathbf{x}_{0})\}$ of dimension $p\times p$ extracted using binary masks $\textbf{P}_{i}$ at random locations $i$ from the full size pre-processed images $\mathbf{x}_{c}$ and $\mathbf{x}_{0}$, respectively. Consequently, the latent noisy data $\mathbf{x}_{t}^{(i)}$ is also treated as a patch. The low-quality TSMI patches $\mathbf{x}_{c}^{(i)}$ are channel-concatenated to $\mathbf{x}_{t}^{(i)}$ as inputs of the denoising network $\epsilon_{\theta}$ to guide the conditional diffusion process. To improve generalization, data augmentation was employed with random vertical/horizontal spatial flips of image patches at each training iteration. We use $T=1000$ diffusion steps with an evenly spaced linear noise schedule $\beta_{t}$ where $\beta_{0}=0.0001$ and $\beta_{T}=0.02$. Our base model uses patch size $p=64$, which we experimentally show that is a good compromise between accuracy and computational time. Training optimizes the loss function in (6) using ADAM optimiser for 150k iterations with fixed learning rate $10^{-4}$, effective batch size 32, dropout rate $0.3$, and an Exponential Moving Average scheme with exponential rate 0.9999 to average model parameters for stable training [21].
@@
6:     $\epsilon_{t}\sim\mathcal{N}(\textbf{0},\textbf{Id})$
@@
8:           $\nabla_{\theta}||\epsilon_{t}-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}^{(i)}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t},\mathbf{x}_{c}^{(i)},t)||^{2}+\lambda L_{\sigma}$
@@
3.3 Iterative Reconstruction by Reverse Diffusion
The reverse diffusion process for MRF reconstruction, that is obtaining restored TSMI samples $\hat{\mathbf{x}}_{0}$ from an initial gridding reconstruction as the conditioning image $\mathbf{x}_{c}$, is detailed in Algorithm 2. We divide the TSMI image into overlapping spatial patches. Staring with each patch $\mathbf{x}_{T}^{(i)}$ initialized as pure noise, the network $\epsilon_{\theta}$ performs reverse iterations, where it takes the conditioned patch $\mathbf{x}_{c}^{(i)}$ alongside the noisy patch $\mathbf{x}_{t}^{(i)}$ to denoise the latter and estimate $\mathbf{x}_{t-1}^{(i)}$. Similar to training, each conditional image is range-normalized to [-1, 1]. Our strategy for patch inference consists of running the complete reverse process on each patch, using IDDPM as a fast sampling backbone, which allows for reconstruction in as few as $K=50\ll T$ steps. After retrieving all overlapping patches of the restored image, they are aggregated and averaged in the common regions (lines 11 to 14). As shown in our experiments, careful selection of patch size and stride during the reverse process is essential for balancing accuracy and computational efficiency. Finally, the reconstructed full-size TSMI, $\hat{\mathbf{x}}_{0}$, is processed by dictionary matching to obtain the T1 and T2 parameter maps. Repeating Algorithm 2 with different random draws for the initial noise $\mathbf{x}_{T}$ enables sampling variation across generated reconstructions, for a given condition image $\mathbf{x}_{c}$.
@@ 3.3 Iterative Reconstruction by Reverse Diffusion
10:         $\mathbf{x}_{t_{\text{next}}}^{(i)}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(\mathbf{x}_{c}^{(i)},\mathbf{x}_{t},t)\right)+{\boldsymbol{\sigma}}_{t}\textbf{z}$
@@ 4.1 Experimental Setup
The proposed method was evaluated along with the baseline approaches on 2D healthy volunteer brain MRF data, acquired using the Steady State Precession (FISP) with the same flip angle schedule as in [3] with repetition time, echo time and inversion time (TR/TE/TI) of 10/1.908/18 ms, respectively, $l=1000$ repetitions (timeframes), non-Cartesian k-space sampling with variable density spiral readouts, matrix size of $n=230\times 230$ with 1mm in-plane resolution and 5mm slice thickness. Data was acquired on a 3T MRI scanner (MR750w system GE HealthCare, Waukesha, WI) with an 8-channel receive-only head RF coil. Our described pipeline was employed for the reconstruction of $R=5$ fold accelerated scans by retrospective truncation of the FISP sequence length to $l=200$. A total of 8 subjects, 15 axial slices each were available. From this, we use 6 subjects for training and the rest for evaluation/testing. The reference Q-Maps were reconstructed using the baseline LRTV method [40] from full-length ($l=1000$) FISP acquisitions. To evaluate the performance of TSMI reconstructions, we employed references estimated from the clean Q-Maps following the Bloch response model defined in Eq. (1).
Tested algorithms, except SCQ, used dictionary matching to map the reconstructed TSMIs to T1 and T2 maps. For this purpose, an MRF dictionary consisting of 95k fingerprints was simulated. The dictionary was created using a logarithmic grid with $(T1,T2)\in[0.01,6]\text{s}\times[0.004,4]\text{s}$) using the Extended-Phase-Graph formalism  [41]. All experiments used a small NVIDIA 2080Ti GPU for computations, and for the training and testing of all deep learning pipelines.

### v2

# Denoising Diffusion Probabilistic Models for Magnetic Resonance Fingerprinting

@@ 1 Introduction
Quantifying the intrinsic nuclear magnetic resonance (NMR) properties of tissues [1], such as T1 and T2 relaxation times, provides a powerful tool for tissue characterization and monitoring of pathological changes. However, conventional quantitative MRI (qMRI) techniques face challenges due to long acquisition times, limiting their clinical utility. This is because qMRI often requires sampling a long sequence (time series) of magnetization images to capture a single tissue property per scan, making it difficult to acquire the full spatiotemporal volume required for encoding the parameter information of interest in a reasonable scan time. As encoding-efficient alternatives to conventional steady-state parameter mapping techniques, Magnetic Resonance Fingerprinting (MRF) [2] and other transient-state imaging techniques [3, 4, 5] were introduced to address this problem. To reduce scan time, MRF uses short-length acquisition sequences to encode multiple tissue properties simultaneously and applies compressed sensing to subsample only a fraction of the spatiotemporal k-space data. However, faster scans lead to challenges in image reconstruction, including aliasing artifacts from k-space undersampling and limited tissue property information due to the truncated acquisition sequences. Effective image reconstruction algorithms are needed to tackle these challenges and improve the accuracy and precision of tissue parameter estimation.
MRF Imaging Algorithms
@@
Early MRF works [2, 6] used gridding reconstruction (non-Cartesian Fourier back-projection) to convert undersampled k-space data into image time series. Tissue properties were then estimated through dictionary matching, where voxel time signals were compared to a precomputed dictionary of theoretical tissue fingerprints from Bloch equations. However, the accuracy of these methods was limited by aliasing artifacts that were not addressed during the image reconstruction step. Later compressed sensing models used iterative dictionary matching [7, 8, 9] to enforce time-domain priors during reconstruction. Papers [7, 10] showed that a low-rank subspace model driven by the Singular Value Decomposition (SVD) of MRF dictionary could efficiently act as a time-domain prior, while reducing significant computation overhead. Subsequent compressed sensing approaches [11, 12, 13, 14] additionally incorporated some form of sparsity and/or low rank spatial priors to reduce undersampling artefacts. However, the most effective MRF algorithms to date are data driven, based on deep learning. The initial MRF deep learning works [15, 16, 17] replaced the dictionary matching step with pixel-wise neural network models trained solely on the MRF dictionary, without incorporating anatomical information. Consequently, these models had limited capacity to handle spatially correlated aliasing artifacts. More recent deep learning works [18, 19, 20, 21, 22, 23] use anatomical MRF imaging datasets to train CNNs with various architectures and losses to learn more effective spatiotemporal priors directly from data. Some models e.g. [21] aim to restore (de-alias) MRF image time series, while tissue quantification is performed by standard dictionary matching. Other models e.g. [19] combine the two steps and directly output restored quantitative maps. Reconstructing MRF image time series, as opposed to directly estimating parameter maps, could provide a more transparent pipeline, ensuring parameter maps adhere to Bloch equations and, by extension, the underlying physics of the model.
Diffusion models
@@
Denoising diffusion probabilistic models (DDPM) [24, 25, 26, 27] have recently gained significant attention due to their remarkable performance on image generation [28, 29, 30] and image restoration tasks [31, 32, 33, 34, 35]. Built on a hierarchy of denoising autoencoders, DDPMs are trained to iteratively refine image details by denoising and reversing a diffusion process. By incorporating conditional information, DDPMs can be guided during sample image generation [26], for example, in image restoration tasks, the model is conditioned on low-quality degraded image inputs, producing high-quality restored images as the outcome [34, 33]. While DDPM applications in medical imaging are emerging (e.g. in segmentation [36, 37], image-to-image translation [38, 39], and image reconstruction [40, 41, 42, 43, 44]), their potential for MRF is yet unexplored. Compared to CNNs, the current state-of-the-art for MRF reconstruction, DDPMs offer several advantages. DDPMs’ iterative refinement allows for stage-wise, progressive restoration of image details from coarse to fine scales, leading to more precise recovery compared to the single-pass structure of CNNs. They also incorporate advanced mechanisms like self-attention [45] or transformers [46], which more effectively model long-range pixel dependencies. Additionally, while CNNs treat image restoration as a deterministic mapping from a corrupted to a restored image, DDPMs operate within a probabilistic framework, allowing them to learn the distribution of possible solutions for a given task and generate multiple samples i.e. a useful feature for capturing uncertainties in the reconstructed images.
Our work
@@
This paper presents, to the best of our knowledge, the first diffusion-based model for MRF reconstruction to enable efficient multi-parametric estimations of T1 and T2 relaxation maps from highly accelerated acquisitions. Our approach, coined as MRF-IDDPM, is trained as a conditional DDPM model to generate restored MRF image time series from initial gridding reconstructions, which suffer from significant aliasing artifacts due to accelerated MRF acquisitions. Evaluated on in-vivo brain scans data, we show that the proposed approach outperforms established deep learning and compressed sensing MRF algorithms in terms of image reconstruction and quantitative mapping accuracies. Additionally, the probabilistic nature of DDPM allows us to generate multiple sample reconstructions given an input and produce a variance map to visualize reconstruction uncertainties. To address the computational challenges associated with training diffusion models on high-dimensional MRF data, we employed a subspace dimensionality reduction approach from [6] to temporally compress the MRF image time series, enabling the diffusion process to operate within a reduced latent space. We further implemented a patch-based processing pipeline to efficiently handle the spatial dimensions of data, enabling training and reconstruction on modest GPUs with limited memory and reasonable runtimes.
The rest of the manuscript is structured as follows: in Section 2 we provide the preliminaries for our method, which is then described in more details in Section 3. Numerical evaluations and comparisons with baselines techniques are provided in Section 4, alongside the results of ablation studies and a discussion of limitations and future directions. Our conclusions are provided in Section 5.
@@ 2.1 The MRF Reconstruction Problem
The MRF reconstruction is an inverse problem:
@@
[…]
The MRF acquisition process can be greatly shortened by combining k-space undersampling with a reduction in the number of measured time frames $l$. However this approach complicates the reconstruction problem by introducing significant aliasing artifacts, limited tissue parameter encoding and increased noise sensitivity — challenges we aim to address in this work.
@@ 2.2 Denoising Diffusion Probabilistic Models for Image Reconstruction
Denoising Diffusion Probabilistic Models (DDPM) [24, 25] describe a Markovian process in which a sample image from the data distribution $q(\mathbf{x}_{0})$ is gradually corrupted by adding Gaussian noise over $t=1,\ldots,T$ diffusion steps, generating latent variables $\mathbf{x}_{t}$ with increasing noise levels $\beta_{t}\in(0,1)$. For a sufficiently large $T$, the variable $\mathbf{x}_{T}$ approximates pure noise with an isotropic Gaussian distribution. The stepwise forward process is defined as:
$q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\textbf{Id}),$ (2)
@@
without having to progress through all previous diffusion steps. DDPMs train a neural network $\epsilon_{\theta}$ to estimate the noise $\epsilon$ at each diffusion step, with a learning objective to minimize the loss:
@@ 3.1 Preprocessing
In the preprocessing step, we start with k-space measurements $\mathbf{y}_{\text{ref}}$ from extended reference MRF acquisitions such as [3]. These can be used to generate low-aliasing quantitative maps via compressed sensing reconstruction methods like LRTV [47]. From these quantitative maps we then produce high-quality (low-aliasing) reference TSMIs $\mathbf{x}_{\text{ref}}$ using the Bloch response model in Eq. (1). Next, we subsample $\mathbf{y}_{\text{ref}}$ by truncating it along the time dimension to obtain k-space data $\mathbf{y}$, representing retrospectively faster acquisitions. Reconstruction of these accelerated acquisitions using a basic gridding algorithm $\mathbf{x}_{\text{grid}}=A^{H}\mathbf{y}$, or a compressed sensing method leads to severe aliasing artifacts.
@@ 3.2 Training
The training process is outlined in Algorithm 1. We choose a denoising model $\epsilon_{\theta}$ from [26]. The model is trained on TSMI patches $\{\mathbf{x}_{c}^{(i)}=\text{Crop}(\textbf{P}_{i},\mathbf{x}_{c})$, $\mathbf{x}_{0}^{(i)}=\text{Crop}(\textbf{P}_{i},\mathbf{x}_{0})\}$ of dimension $p\times p$ extracted using binary masks $\textbf{P}_{i}$ at random locations $i$ from the full size pre-processed images $\mathbf{x}_{c}$ and $\mathbf{x}_{0}$, respectively. Consequently, the latent noisy data $\mathbf{x}_{t}^{(i)}$ is also treated as a patch. The low-quality TSMI patches $\mathbf{x}_{c}^{(i)}$ are channel-concatenated to $\mathbf{x}_{t}^{(i)}$ as inputs of the denoising network $\epsilon_{\theta}$ to guide the conditional diffusion process. To improve generalization, data augmentation was employed with random vertical/horizontal spatial flips of image patches at each training iteration. We use $T=1000$ diffusion steps with an evenly spaced linear noise schedule $\beta_{t}$ where $\beta_{0}=0.0001$ and $\beta_{T}=0.02$. Our base model uses patch size $p=64$, which we experimentally show to be a good compromise between accuracy and computational time. Training optimizes the loss function in (6) using ADAM optimiser for 150K iterations with fixed learning rate $10^{-4}$, effective batch size 32, dropout rate $0.3$, and an exponential moving average scheme with exponential rate 0.9999 to average model parameters for stable training [26].
@@
6:     $\epsilon\sim\mathcal{N}(\textbf{0},\textbf{Id})$
@@
8:           $\nabla_{\theta}||\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}^{(i)}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t,\mathbf{x}_{c}^{(i)})||^{2}+\lambda L_{\sigma}$
@@
3.3 Inference by Reverse Diffusion
Algorithm 2 describes the iterative reverse diffusion process for MRF reconstruction so to obtain restored TSMI samples $\hat{\mathbf{x}}_{0}$ from an initial gridding reconstruction as the conditioning image $\mathbf{x}_{c}$. We divide the TSMI image into overlapping spatial patches. Starting with each patch $\mathbf{x}_{T}^{(i)}$ initialized as pure noise, the network $\epsilon_{\theta}$ performs reverse iterations, where it takes the conditioned patch $\mathbf{x}_{c}^{(i)}$ alongside the noisy patch $\mathbf{x}_{t}^{(i)}$ to denoise the latter and estimate $\mathbf{x}_{t-1}^{(i)}$. Similar to training, each conditional image is range-normalized to [-1, 1]. Our strategy for patch inference consists of running the complete reverse process on each patch, using IDDPM as a fast sampling backbone, which allows for reconstruction in as few as $K=50\ll T$ steps. After retrieving all overlapping patches of the restored image, they are aggregated and averaged in the common regions (lines 11 to 14). As shown in our experiments, careful selection of patch size and stride during the reverse process is essential for balancing accuracy and computational efficiency. Finally, the reconstructed full-size TSMI, $\hat{\mathbf{x}}_{0}$, is processed by dictionary matching to obtain the T1 and T2 parameter maps. Repeating Algorithm 2 with different random draws for the initial noise $\mathbf{x}_{T}$ enables sampling variation across generated reconstructions, for a given condition image $\mathbf{x}_{c}$.
@@ 3.3 Inference by Reverse Diffusion
10:         $\mathbf{x}_{t_{\text{next}}}^{(i)}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}^{(i)}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(\mathbf{x}_{t}^{(i)},t,\mathbf{x}_{c}^{(i)})\right)+{\boldsymbol{\sigma}}_{t}\textbf{z}$
@@ 4.1 Experimental Setup
The proposed method was evaluated along with the baseline approaches on 2D healthy volunteer brain MRF data, acquired using the Steady State Precession (FISP) with the same flip angle schedule as in [3] with repetition time, echo time and inversion time (TR/TE/TI) of 10/1.908/18 ms, respectively, $l=1000$ repetitions (time frames), non-Cartesian k-space sampling with variable density spiral readouts, matrix size of $n=230\times 230$ with 1mm in-plane resolution and 5mm slice thickness. Data was acquired on a 3T MRI scanner (MR750w system GE HealthCare, Waukesha, WI) with an 8-channel receive-only head RF coil.  [^1] This work was conducted under institutional approval for EPSRC grant EP/X001091/1. The retrospective study used anonymized human scans provided by GE Healthcare, with informed consent obtained in compliance with the German Medical Devices Act. Our described pipeline was employed for the reconstruction of $R=5$ fold accelerated scans by retrospective truncation of the FISP sequence length to $l=200$. A total of 8 subjects, 15 axial slices each were available. From this, we use 6 subjects for training and the rest for evaluation/testing. The reference Q-Maps were reconstructed using the baseline LRTV method [47] from full-length ($l=1000$) FISP acquisitions. To evaluate the performance of TSMI reconstructions, we employed references estimated from the clean Q-Maps following the Bloch response model defined in Eq. (1). All experiments used a small NVIDIA 2080Ti GPU for computations, and for the training and testing of all deep learning pipelines.
Method MAPE (%) $\downarrow$ RMSE (ms) $\downarrow$ NRMSE $\downarrow$ SSIM $\uparrow$
T1 T2 T1 T2 T1 T2 TSMI T1 T2 TSMI
SCQ 9.58 23.67 16.22 9.16 0.1143 0.4354 0.9750 0.9674
MRF-IDDPM (Ours) 7.19 17.64 14.10 7.63 0.1000 0.3669 0.4142 0.9759 0.9712 0.8729
Table 1: Reconstruction metrics (averaged over test dataset) for TSMIs and T1/T2 parameter maps using evaluated algorithms.